I need help with this problem plzz!!
Solve the following equation exactly: log(x+1)+log2=1

Answer by mananth(16949)

log(x+1)+log2=1
log(x+1)+log2=log10
log(2(x+1))=log10
2(x+1)=10
2x+2=10
2x=8
x=4